# Portable Targeted Sampling Framework Using LLVM

### v1

@@ Abstract
Comprehensive architectural evaluation of full workloads is throttled by slow simulation and per-binary sampling pipelines. We present Nugget, a flexible framework for portable sampling across simulators and real hardware, ISAs, and libraries. Nugget operates at the LLVM IR level to perform binary-agnostic interval analysis, then emits lightweight, cross-platform executables—nuggets—that can be validated on real machines before driving simulation. Across SPEC CPU2017, NPB, and LSMS, Nugget cuts interval-analysis cost by orders of magnitude relative to functional simulation (up to $\approx 578\times$ on multithreaded NPB), keeps single-thread overhead low, and enables native-speed validation of selected samples. Case studies with gem5 show that nuggets support evaluation of system performance and model accuracy. Nugget makes sampling methodology research faster and more portable.
@@ I Introduction
To address these limitations, we present a framework called Nugget for implementing diverse sample selection methodologies. Because modern applications exhibit varied characteristics, we believe there is no one-size-fits-all sampling methodology or selection algorithm [13, 17, 36, 21]. Therefore, Nugget is designed to be flexible and extensible, allowing researchers to create new sample selection algorithms that can be tailored to specific applications and workloads without worrying about the current limitations of targeted sampling techniques. Additionally, using Nugget for sample selection methodologies, because of its nature of cross-platform compatibility and its ability to run on real hardware, enables richer insights into the selected samples and the systems that run the samples.
The main contributions of this paper are:
@@
  • A LLVM IR level unit of work, enabling sample selection algorithms to create and locate samples for a program in a cross-platform manner, agnostic of machine level instructions, such as architectural specific instructions and optimizations.
@@
  • An efficient validation methodology for validating the selected samples for the specific workload in native time.
@@ II-A Samples Are Expensive To Find
For example, the SPECspeed suite in the SPEC2017 benchmark suite [10] has an average of 21.8 trillion dynamic instructions per benchmark for floating point benchmarks with reference input size, which is the only input size designed to report time metrics [23]. Functional simulators, such as the gem5 atomic model, are often limited to a simulation rate of approximately 1 Million Instructions Per Second (MIPS) [5]. Therefore, it takes approximately 250 days of wallclock time to analyze an average SPECspeed benchmark with reference input size. More recent workloads (e.g., the forthcoming SPEC CPUv8 [33], HPC workloads, and machine learning workloads) are significantly larger, making functional simulation infeasible.
@@ II-B Samples Are Tied To A Single Executable Binary
Most prior work relies on binary-specific information, such as the program counter, to define intervals, identify program phases, and select samples. Any binary change, such as recompiling with different ISA extensions or compilers, alters the static instructions and the dynamic instruction stream, invalidating previously selected samples. As discussed in Section II-A, the high overhead of this analysis makes repeating it for every binary revision time-consuming..
@@ II-C Validating The Selected Samples Is Infeasible
Sampling methodologies aim to predict overall program performance by measuring only a subset of representative samples and extrapolating the full-program behavior using statistical models. Most methodologies focus on predicting metrics that summarizes the entire hardware performance, such as instructions per cycle (IPC) or total runtime.
@@
The effectiveness of the sample selection method can vary based on the benchmark’s characteristics. Thus, the selected representative intervals must be validated to ensure that they are representative of the entire workload. However, the only way to validate the representative intervals is to run the entire workload in detailed simulation, which is infeasible for realistic workloads.
@@
This limitation has led to sampling techniques being rarely validated with large benchmarks. There is no guarantee that the validation results of smaller benchmarks or partial executions of benchmarks can be generalized to large benchmarks. Moreover, because it is difficult to obtain the actual program phases of the workloads, sampling methodologies typically have little insight into why one sampling technique works better than another, or why not. Instead, they often only compare the final predicted metric results on particular sets of benchmarks. This can be misleading, either by chance or by exploiting the effects of having a large enough sample size.
Goal 3: Quickly validate the selected samples for the targeted benchmarks and input size.
@@ III The Nugget Framework
Figure 1: Nugget pipeline. The process is divided into two main stages: (1) preparation and analysis, and (2) nugget selection and validation. The preparation and analysis stage, including base IR file creation and interval analysis execution, only needs to be performed once per program. In contrast, nugget creation and validation can be repeated to refine and select the most representative set of samples for the workload.
Nugget is designed to address the limitation of the prior sampling techniques, specifically focusing on 1) the high cost of finding samples, 2) the coupling of the samples to a specific executable binary, and 3) the infeasibility of validating the selected samples. Nugget leverages LLVM passes to statically insert annotations to analyze the program’s execution on real hardware without the need for simulation, emulation, or binary translation. By using LLVM intermediate representation (IR) to define program progress and intervals, the samples become independent of the binary and can be reused across different executable versions of the same program, since LLVM IR abstracts away details of the final binary. Lastly, Nugget provides a method to create portable samples to be executed on any platform, such as real hardware or simulators, so that the samples can be validated in native time with real hardware before running them on a simulator.
@@
We begin by defining program progress and intervals, which ensures samples maintain independent of the binary in Section III-A. Next, we describe the preparation required by Nugget to guarantee the LLVM IR of the program remain consistent across different binaries in Section III-B. Then, we introduce our novel interval analysis methodology, which leverage LLVM passes, enabling comprehensive program analysis directly on real hardware without imposing hardware-specific constraints in Section III-C. Following that, we present the nugget creation process, which allows for the generation of portable samples that can be executed on any platform, including real hardware and simulators, in Section III-D. Finally, we describe the sample selection validation process, which enables efficient validation of the selected samples for specific workloads in native time, ensuring that the samples accurately represent the program’s execution in Section III-E.
@@ III-B Preparation
Optimization is important for the program’s performance. LLVM IR level optimizations, such as loop unrolling, can significantly affect control flow. To maintain consistency, Nugget applies these optimizations during the generation of the base LLVM IR representation in the preparation stage. In contrast, most backend optimizations, such as replacing scalar instructions with AVX vector instructions, do not alter the structure of the LLVM IR. While certain backend passes (e.g., late-stage loop unrolling) may modify LLVM IR layout under specific conditions, such passes are not enabled by default in LLVM. Therefore, we can apply the backend optimization in the final binary compilation safely.
@@ III-D1 Markers
Tracking every executed IR instruction incurs prohibitive overhead. To avoid this, we designate the last executed LLVM IR basic block (IRBB) that completes the target instruction count as a marker. This marker consists of two components: the IRBB itself and the number of times it must execute. When this IRBB executes for the final required count, it implies that the total desired number of IR instructions has been reached, due to the control flow properties of execution. By tracking only executions of selected IRBBs, we achieve nearly the same precision as full IR-level tracking and substantially reduced overhead.
Nugget enables cross-platform execution by inserting hooks at these marker IRBBs through an LLVM pass. These hooks track IRBB executions and trigger specific actions, such as resetting statistics in a simulator or starting a timer on real hardware. The end marker for an interval is determined using the IRBB vector and count-stamp vector gathered during interval analysis (Section III-C2). Specifically, we use the count-stamp vector (Section III-C2) to identify the last executed IRBB of the interval, which becomes the basic block for the end marker, and the IRBB vector to determine how many times that block must execute to reach the interval’s endpoint. Because the start of one interval corresponds to the end of the previous interval, we reuse the end marker of the preceding interval as the start marker for the next.
When using nuggets in simulation, we can also mark the start of a warmup interval. The warmup interval does not belong to the selected sample but is a period of execution used to warm up the microarchitectural state in simulation models, assuming functional simulation is used to fast-forward execution before the sample starts [15].
@@ III-E Sample Validating
To run a nugget on real hardware, the system must still execute the portion of the program leading up to the sample’s start marker (Section III-D1). For samples that are in later stages of a large program’s execution, this can introduce some validation overhead. However, this cost is minor compared to the substantial overhead of simulating the entire program.
In the future, we plan to reduce this overhead further by exploring techniques such as creating software checkpoints using CRIU [6] to resume execution closer to the sample start point.
@@ IV Evaluation
    Quickly validate the selected samples for the targeted benchmarks and input size.
@@
Table I presents the detailed configurations of the real hardware used in our evaluation. For the simulation models, we employ a basic generic functional model using the ATOMIC CPU in gem5, running a guest Ubuntu 24.04 environment, to perform interval analysis on both x86 and Arm architectures.
@@
We evaluate Nugget using the SPEC2017 benchmark suite, the NPB benchmark suite, and the LSMS benchmark, covering a wide range of realistic workloads as summarized in Table II. The input chosen for each benchmark is too large to be simulated in feasible time (except for NPB class “A”). For SPEC2017, we use the reference input sizes, which require between 198 billion and 9,897 billion instructions. For NPB, we use class “C” and “D” inputs, with class “C” requiring between 24 billion and 2,714 billion instructions and class “D” ranging from 1,217 billion to 37,533 billion instructions. For LSMS, we use the Fe input, which requires between 28,484 billion and 198,885 billion instructions (one binary per ISA).
@@ IV-A1 Comparison with Functional Simulation
For this experiment, we use the NPB benchmark suite with Class “A” inputs and 4 threads. We select Class “A” because larger input sizes, such as those used in Figure 3, would require months to analyze using functional simulation. We use multi-threaded workloads to demonstrate a worst-case overhead scenario, since interval analysis must track synchronization events between threads. Additionally, multi-threaded workloads reflect current trends in modern computing, making them a relevant evaluation target.
@@
Overall, multi-threaded workloads amplify interval analysis overhead because their baseline execution times are significantly shorter—due to parallelism—while the analysis itself incurs greater cost from synchronization requirements. The single-threaded workloads in Figure 3 exhibit much lower overheads, which we discuss further in Section IV-A2. The “cg” benchmark has the shortest execution time in this experiment (0.28 seconds), which further amplifies the measured overhead for both methods. The “is” benchmark exhibits the lowest overhead under functional simulation but still incurs a $95\times$ slowdown with Nugget, primarily due to the frequent executions of interval analysis hooks, as discussed in Section IV-B3. The “mg” and “ft” benchmarks show relatively high overhead for both methods because their parallel regions increase synchronization costs during interval analysis, which in turn adds further simulation overhead. We examine the relationship between synchronization and overhead in more detail in Section IV-A3.
@@ IV-A2 Mix of Workload Types
We also evaluate overhead across a mix of workload types. For single-threaded workloads, we use the SPEC 2017 suite with the ref input size and the LSMS benchmark with the Fe input size. For multi-threaded workloads, we use the NPB benchmark suite with Class “C” in this experiment.
Figure 3 shows the interval analysis slowdown across the evaluated workloads and demonstrates that the overhead varies significantly depending on the level of synchronization. The average observed overhead is $3\times$ for SPEC 2017, $2\times$ for LSMS, and $34\times$ for NPB. The overhead remains low for SPEC 2017 and LSMS because they are single-threaded workloads, which simplifies interval analysis by eliminating the need for synchronization. In contrast, the high overhead for NPB arises from the synchronization required during multi-threaded interval analysis, particularly due to aggregation across threads. As synchronization demands increase, so does the overhead, a relationship we explore further in Section IV-A3.
To demonstrate the flexibility of our approach across host machines, we repeat the interval analysis on two distinct systems listed in Table I and find consistent results across both platforms for single-threaded workloads, with one corner case. For SPEC 2017, interval analysis results are identical across the two systems. For NPB, the multi-threaded nature of the benchmarks introduces variability between runs (even on the same host), a well-known challenge in multi-threaded sampling that is beyond the scope of this work [30, 12]. For LSMS, we observe minor differences in the execution frequency of some IRBBs, which we attribute to differences in floating-point precision across CPUs. In particular, the number of iterations in certain data-driven loops in LSMS is influenced by the underlying architecture’s precision. Such differences could affect nugget representativeness if a marker IRBB is sensitive to these variations. Nevertheless, as shown in Section IV-B2, nuggets generated from interval analysis on one system still achieve low prediction error when validated on the other system for LSMS.
@@ IV-A3 Scaling with Number of Threads
Finally, we study how interval analysis overhead scales with the number of threads. For this experiment, we use the NPB benchmark suite with Class “D” inputs and vary the thread count across {1, 4, 8, 16}. To ensure fair comparisons, all overheads are normalized to the execution time of the original workload running with a single thread.

### v2

# Nugget: Portable Program Snippets

@@ Abstract
Evaluating architectural ideas on realistic workloads is increasingly challenging due to the prohibitive cost of detailed simulation and the lack of portable sampling tools. Existing targeted sampling techniques are often tied to specific binaries, incur significant overhead, and make rapid validation across systems infeasible. To address these limitations, we introduce Nugget, a flexible framework that enables portable sampling across simulators, hardware, architectural differences, and libraries. Nugget leverages LLVM IR to perform binary-independent interval analysis, then generates lightweight, cross-platform executable snippets, nuggets, that can be validated natively on real hardware before use in simulation. This approach decouples samples from specific binaries, dramatically reduces analysis overhead, and allows researchers to iterate on sampling methodologies while efficiently validating samples across diverse systems.
@@ I Introduction
To address these limitations, we present a framework called Nugget for agile targeted sampling development. Because modern applications exhibit varied characteristics, we believe there is no one-size-fits-all sampling methodology or selection algorithm [16, 24, 51, 28]. Therefore, Nugget is designed to be flexible and extensible, allowing researchers to create new sample selection algorithms that can be tailored to their targeted workload and input without worrying about the current limitations of targeted sampling techniques. Additionally, using Nugget for sample selection methodologies, because of its nature of cross-platform compatibility and its ability to run on real hardware, enables richer insights into the selected samples and the systems that run the samples. The main contributions of this paper are:
@@
  • An LLVM IR level unit of work, enabling sample selection algorithms to create and locate samples for a program in a cross-platform manner, agnostic of machine level instructions, such as architectural specific instructions and optimizations.
@@
  • An efficient validation methodology that validates the selected samples for the target workload and input on native hardware, integrated into the development workflow.
@@ II-A Samples Are Expensive To Find
For example, the SPECspeed suite in the SPEC CPU2017 benchmark suite [11] has an average of 21.8 trillion dynamic instructions per benchmark for floating point benchmarks with reference input size, which is the only input size designed to report time metrics [31]. Functional simulators, such as the gem5 atomic model, are often limited to a simulation rate of approximately 1 Million Instructions Per Second (MIPS) [5]. Therefore, it takes approximately 250 days of wall-clock time to analyze an average SPECspeed benchmark with reference input size. More recent workloads (e.g., the forthcoming SPEC CPUv8 [44], HPC workloads, and machine learning workloads) are significantly larger, making functional simulation infeasible.
@@ II-B Samples Are Tied To A Single Executable Binary
Most prior work relies on binary-specific information, such as the program counter, to define intervals, identify program phases, and select samples. Any binary change, such as recompiling with different ISA extensions or compilers, alters the static instructions and the dynamic instruction stream, invalidating previously selected samples. As discussed in Section II-A, the high overhead of this analysis makes repeating it for every binary revision time-consuming.
@@ II-C Validating The Selected Samples Is Infeasible
Sampling methodologies aim to predict overall program performance by measuring only a subset of representative samples and extrapolating the full-program behavior using statistical models. Most methodologies focus on predicting metrics that summarize the entire hardware performance, such as instructions per cycle (IPC) or total runtime.
@@
The effectiveness of a sample selection method depends on benchmark characteristics. Differences in control flow phase behavior, working set size and locality, and input sensitivity can cause a method that succeeds on one workload and input pair to misrepresent another [51, 41, 48, 15, 16, 19, 12]. For example, empirical studies (e.g., MAV [12]) report outliers where SimPoint fails to capture distinct memory access behavior while SimPoint reports average 3% IPC error with SPEC CPU2000 [42] programs [26]. Hence, representative intervals should be validated. However, the only way to validate them is to run the entire workload.
@@
Lack of an agile workflow to validate quickly has led to sampling techniques being rarely validated with large benchmarks. Evidence suggests that validation results from smaller benchmarks or partial executions do not generalize to larger benchmarks and realistic workloads [51, 48, 41, 31, 26]. Moreover, the program behavior of a workload with different input can be different [15]. Thus, per benchmark and input pair validation is needed.
Goal 3: Quickly validating the selected samples for the targeted benchmarks and input size.
@@ III The Nugget Framework
Nugget is designed to address the limitations of the prior sampling techniques, specifically focusing on 1) the high cost of finding samples, 2) the coupling of the samples to a specific executable binary, and 3) the infeasibility of validating the selected samples. Nugget leverages LLVM passes to statically insert annotations to analyze the program’s execution on real hardware without the need for simulation, emulation, or binary translation. By using LLVM intermediate representation (IR) to define program progress and intervals, the samples become independent of the binary and can be reused across different executable versions of the same program, since LLVM IR abstracts away details of the final binary. Lastly, Nugget provides a method to create portable samples to be executed on any platform, such as real hardware or simulators, so that the samples can be validated in native time with real hardware before running them on a simulator.
@@
We begin by defining program progress and intervals, which ensures that samples remain independent of the binary, as described in Section III-A. Next, we describe the preparation required by Nugget to guarantee the LLVM IR of the program remains consistent across different binaries in Section III-B. Then, we introduce our novel interval analysis methodology, which leverages LLVM passes, enabling comprehensive program analysis directly on real hardware without imposing hardware-specific constraints in Section III-C. Following that, we present the nugget creation process, which allows for the generation of portable samples that can be executed on any platform, including real hardware and simulators, in Section III-D. Finally, we describe the sample selection validation process, which enables efficient validation of the selected samples for specific workloads in native time, ensuring that the samples accurately represent the program’s execution in Section III-E.
Figure 1: Nugget pipeline. The process is divided into two main stages: (1) preparation and analysis, and (2) nugget selection and validation. The preparation and analysis stage, including base IR file creation and interval analysis execution, only needs to be performed once per program. Nugget creation and validation can be repeated to refine and select the most representative set of samples for the workload.
@@ III-A Unit of Work
Therefore, Nugget’s unit of work is suitable for applications that have stable control flow across different platforms. Applications such as gradient descent, where floating-point precision might affect the control flow, are not suitable for Nugget or for any sampling methodologies that rely on the application having stable control flows across microarchitectures [26, 41, 39, 36, 32, 14, 50, 30].
@@ III-B Preparation
Optimization is important for the program’s performance. LLVM IR level optimizations, such as loop unrolling, can significantly affect control flow. To maintain consistency, Nugget applies these optimizations during the generation of the base LLVM IR representation in the preparation stage. In contrast, most backend optimizations, such as replacing scalar instructions with AVX vector instructions, do not alter the structure of the LLVM IR. While certain backend passes (e.g., late-stage loop unrolling) may modify LLVM IR layout under specific conditions, such passes are not enabled by default in LLVM.
While build-time modes such as LTO and PGO can alter LLVM IR, they are disabled by default. For our experiments, the LLVM IR structure remained stable across optimization levels, CPU-feature flags, and LLVM versions. Therefore, we can apply the backend optimization in the final binary compilation safely.
@@ III-D1 Markers
We designate the last executed LLVM IR basic block (IRBB) that completes the target instruction count as a marker. This marker consists of two components: the IRBB itself and the number of times it must execute. The end marker for an interval is determined using the IRBB vector and count-stamp vector gathered during interval analysis (Section III-C2). Specifically, we use the count-stamp vector (Section III-C2) to identify the last executed IRBB of the interval, which becomes the basic block for the end marker, and the IRBB vector to determine how many times that block must execute to reach the interval’s endpoint. Because the start of one interval corresponds to the end of the previous interval, we reuse the end marker of the preceding interval as the start marker for the next.
Nugget inserts hooks at these marker IRBBs during the nugget creation LLVM pass. These hooks track IRBB executions and trigger specific actions, such as resetting statistics in a simulator or starting a timer on real hardware.
When using nuggets in simulation, we can also mark the start of a warmup interval. The warmup interval does not belong to the selected sample but is a period of execution used to warm up the microarchitectural state in simulation models, assuming functional simulation is used to fast-forward execution before the sample starts [20]. Any contiguous and antecedent intervals of the sample interval can be used as the warmup interval, so one can choose how many intervals to use for warmup.
@@ III-E Sample Validating
To run a nugget on real hardware, the system must still execute the portion of the program leading up to the sample’s start marker (Section III-D1). For samples that are in later stages of a large program’s execution, this can introduce some validation overhead.
In the future, we plan to reduce this overhead further by exploring techniques such as creating software checkpoints using CRIU [7] or BLCR [23] to resume execution closer to the sample start point.
@@ IV Evaluation
    Quickly validating the selected samples for the targeted benchmarks and input size.
@@
Table II presents the detailed configurations of the real hardware used in our evaluation. For the simulation models, we employ a basic generic functional model using the ATOMIC CPU in gem5, running a guest Ubuntu 24.04 environment, to perform interval analysis on both x86 and Arm architectures.
We evaluate Nugget using the SPEC CPU 2017 benchmark suite, the NAS Parallel Benchmarks (NPB) suite [8], and the LSMS benchmark [46], covering a wide range of realistic workloads as summarized in Table II.
We use the “-O2” optimization level for the compilation. The input chosen for each benchmark is too large to be simulated in feasible time (except for NPB class “A”).
For SPEC CPU2017, we use the reference input sizes, which require between 198 billion and 9,897 billion instructions. For NPB, we use class “C” and “D” inputs, with class “C” requiring between 24 billion and 2,714 billion instructions and class “D” ranging from 1,217 billion to 37,533 billion instructions. For LSMS, we use the Fe input, which requires between 28,484 billion and 198,885 billion instructions (one binary per ISA).
We used SPEC CPU2017 Integer suite, except for “619.lbm_s” and “644.nab_s”, which are from the Floating-Point suite. We do not include CPU 2017 benchmarks that failed to compile with the original SPEC CPU2017 GCC build system (6 benchmarks) or encountered Fortran linking issues (5 benchmarks). We did not encounter the same Fortran linking issues with any NPB Fortran benchmarks.
@@ IV-A1 Comparison with Functional Simulation
For this experiment, we use the NPB suite with Class “A” inputs and 4 threads. We select Class “A” because larger input sizes, such as those used in Figure 3, would require months to analyze using functional simulation. We use multi-threaded workloads to demonstrate a worst-case overhead scenario, since interval analysis must track synchronization events between threads. Additionally, multi-threaded workloads reflect current trends in modern computing, making them a relevant evaluation target.
@@
Overall, multi-threaded workloads amplify interval analysis overhead because their baseline execution times are significantly shorter (due to parallelism) while the analysis itself incurs greater cost from synchronization requirements. The single-threaded workloads in Figure 3 exhibit much lower overheads, which we discuss further in Section IV-A2. The “cg” benchmark has the shortest execution time in this experiment (0.28 seconds), which further amplifies the measured overhead for both methods. The “is” benchmark exhibits the lowest overhead under functional simulation but still incurs a $95\times$ slowdown with Nugget, primarily due to the frequent executions of interval analysis hooks, as discussed in Section IV-B3. The “mg” and “ft” benchmarks show relatively high overhead for both methods because their parallel regions increase synchronization costs during interval analysis, which in turn adds further simulation overhead. We examine the relationship between synchronization and overhead in more detail in Section IV-A3.
@@ IV-A2 Mix of Workload Types
We also evaluate overhead across a mix of workload types. For single-threaded workloads, we use the SPEC CPU2017 suite with the ref input size and the LSMS benchmark with the Fe input size. For multi-threaded workloads, we use the NPB suite with Class “C” in this experiment.
Figure 3 shows the interval analysis slowdown across the evaluated workloads and demonstrates that the overhead varies significantly depending on the level of synchronization. The average observed overhead is $3\times$ for SPEC CPU2017, $2\times$ for LSMS, and $34\times$ for NPB. The overhead remains low for SPEC CPU2017 and LSMS because they are single-threaded workloads, which simplifies interval analysis by eliminating the need for synchronization. In contrast, the high overhead for NPB arises from the synchronization required during multi-threaded interval analysis, particularly due to aggregation across threads. As synchronization demands increase, so does the overhead, a relationship we explore further in Section IV-A3.
To demonstrate the flexibility of our approach across host machines, we repeat the interval analysis on two distinct systems listed in Table II and find consistent results across both platforms for single-threaded workloads, with one corner case. For SPEC CPU2017, interval analysis results are identical across the two systems. For NPB, the multi-threaded nature of the benchmarks introduces variability between runs (even on the same host), a well-known challenge in multi-threaded sampling that is beyond the scope of this work [39, 14]. For LSMS, we observe minor differences in the execution frequency of some IRBBs, which we attribute to differences in floating-point precision across CPUs. In particular, the number of iterations in certain data-driven loops in LSMS is influenced by the underlying architecture’s precision. Such differences could affect nugget representativeness if a marker IRBB is sensitive to these variations. Nevertheless, as shown in Section IV-B2, nuggets generated from interval analysis on one system still achieve low prediction error when validated on the other system for LSMS.
@@ IV-A3 Scaling with Number of Threads
Finally, we study how interval analysis overhead scales with the number of threads. For this experiment, we use the NPB suite with Class “D” inputs and vary the thread count across {1, 4, 8, 16}. To ensure fair comparisons, all overheads are normalized to the execution time of the original workload running with a single thread.